Murray and Max must paint the fence around their home. The fence is 810 feet long. Murray agrees to paint 25% more of the fence than Max. How many feet of fence must each paint?
Answer by mananth(16946)

the ratio of the job they do = Murray : Max :: 1:1.25
multiply by 100
=100:125
(100+125)= 225
..
(100/225)*810
= 360
(125/225)*810
=450
CHECK
(360*1.25)= 450